40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15559     Entry time: Sat Sep 5 14:28:03 2020
Author: Koji 
Type: Update 
Category: General 
Subject: LO beam: Fiber coupling work 

2PM: Arrived at the 40m. Started the work for the coupling of the RF modulated LO beam into a fiber. -> I left the lab at 10:30 PM.

The fiber coupling setup for the phase-modulated beam was made right next to the PSL injection path. (See attachment 1)

  • For the alignment of the beam, the main PSL path, including the alignment of the 2" PO mirror, has not been touched.
  • There are two PO beams with the optical power of 0.8mW (left) and 1.6mW (right). Both had been blocked but the right one was designed to be used for PSL POS and ANG. For the fiber coupling, the right beam was used.
  • The alignment/mode-matching work has been done with a short (2m?) fiber patch cable from Thorlabs. The fiber is the same as the one used for LO delivery.
  • I tried to have a mode-matching telescope in the LO path. I ended up having no lens for the best result. The resulting transmitted power is 1.21mW out of 1.64mW incident (~74%). These powers were measured with the Ophir power meter. (Note that Thorlabs' fiber power meter indicated 1.0mW transmission.)

Some notes

  • After the PSL activity, the IMC locking was checked to see if I messed up the PSL alignment. It locks fine and looks fine.
    • The input shutter (left closed after Jon's vacuum work?) was opened.
    • The alignment was not optimal and had some pitch misalignment (e.g. TEM03).
    • After some MC SUS alignment, the automatic locking of TEM00 was recovered. Mainly MC3 pitch was moved (+0.17).
    • I've consulted with Gautam and he thinks this is with the level of regular drift. The AS beam was visible.
  • The IMC and MI were moving so much, but this seemed just the usual Saturday night Millikan shake.
  • During the activity, the PSL HEPA was turned up to 100 and it was reverted to 33 after the work.
  • I have been wearing a mask and gloves throughout the work there.
Attachment 1: 20200905212254_IMG_9938.JPG  284 kB  Uploaded Sun Sep 6 10:36:09 2020  | Hide | Hide all
20200905212254_IMG_9938.JPG
ELOG V3.1.3-